Sheridan Company has fixed costs of $540000 and variable costs are 40% of sales. How much will Sheridan report as sales when its net income equals $54000?

$990000
$1485000
$954000

$237600

Management of the Vaughn Manufacturing would like the Food Division to transfer 9700 cans of its final product to the Restaurant Division for $27. The Food Division sells the product to customers for $67 per unit. The Food Division’s variable cost per unit is $35 and its fixed cost per unit is $8.

If the Food Division is currently operating at full capacity, what is the minimum transfer price the Food Division should accept?

$43
$67
$27
$35
Answer #1
Solution-1
Correct answer will be $990000
Detail working for refrence
Revised Contribution= Fixed Cost + Net Income
=540000+54000=594000
Contribution Margin Ratio= Sale- variable cost
100-40%=60%
Sales= Revised Contribution Margin/CM Ratio
=$594000/60%=$990000

Solution-2

Minimum Transfer Price Food Division should accept is $67 because food division is currently operated at full operating capacity and product can be sold easily in outside market for $67.